A cable hangs in a parabolic arc between two poles 100 feet apart. The poles are 30 feet high and the lowest point on the suspended cable is 5 feet above the ground. Find the equation of the arc if the vertex is the lowest point of the cable.


Also, find the height of the cable at a point 10 feet from one of the poles.


1
Expert's answer
2021-10-06T14:06:22-0400

Suppose that the equation of the parabola is "y(x)=ax^2+c, a>0."

The lowest point on the suspended cable is 5 feet above the ground


"y(0)=a(0)^2+c=5=>c=5"

The poles are 100 feet apart. The poles are 30 feet high.


"y(-50)=y(50)=30"

Then


"30=a(50)^2+5=>a=\dfrac{35}{2500}=0.014"

a) The equation of the arc is


"y(x)=0.014x^2+5"

b)


"y(40)=0.014(40)^2+5=27.4"




"y(60)=0.014(60)^2+5=55.4"

The height of the cable at a point 10 feet from one of the poles is 27.4 feet or 55.4 feet.